OpenStudy (anonymous):

the velocity of a 0.25kg motorcycle rocket changes from 15m/s [up] to 40m/s [up] in 0.60 seconds. Calculate the force that the escaping gas exerts on the rocket? What formula should i use to calculate this ?

OpenStudy (anonymous):

@experimentX ?

OpenStudy (experimentx):

not sure .. unless i know the answer.

OpenStudy (anonymous):

answer is 13N [up]

OpenStudy (experimentx):

14 or 10 ??

OpenStudy (anonymous):

the book says 13

OpenStudy (anonymous):

what formula do i use?

OpenStudy (experimentx):

Not really sure ... since the rate of change of mass is also not given.

OpenStudy (mos1635):

 for the rocket total Force=M*a F'-W=M*a and F'=F from Newton 3rd low